# What is a good use case for quantum computers?

### v1

@@ I Introduction
Quantum computing hardware has progressed substantially in the last few years. Several quantum computing hardware providers can relatively consistently achieve two qubit gates with more than 99% Fidelity. Under laboratory conditions, even 99.9% has been achieved in quite a few architectures, and at least one cloud-accessible device with these remarkable fidelities does exist (Quantinuum). This all makes it even more important to drive use cases forward. A vast amount of work has been dedicated to testing various algorithms and approaches in the last few years [1, 2, 3, 4, 5, 6] and it is now all the more important to assess the state and quality of the existing ideas for use cases.
@@
However, it is often noticable that combining just these two criteria in a completely satisfactory way remains quite challenging. As we will discuss later, there is a vast number of studies on problems that are without any doubts of great industrial relevance, but lack substantiated arguments on how to achieve quantum advantage. Concurrently, it is highly non-trivial to connect algorithms with strictly provable quantum advantage to actual applications. Most prominently, this is the case for sampling experiments which are certainly of crucial importance as a fundamental performance benchmark, but until recently [7] lacked any connection to applications.
@@
The second criterion we intend to add is the need to invest as much effort to solve the given problem with classical computers as into the research on quantum computing. We call this criteria Get the Job Done Without the Quantum Computer. Of course, scientific research often has limited ressources, so it might not always be possible to fully optimize classical methods before exploring the quantum ones. Nevertheless, at some stage in the process, it’s important to consider the status and availability of optimized classical solutions when judging any work on a quantum use case. As an example, one of the first claims advanced for an andvantage in using quantum computers was based on a benchmark performed by a D-Wave quantum computer [8] versus classical CPLEX solver, a standard tool for combinatorial optimization problems. However, this example did not stand up to scrutiny in the long run, since CPLEX is not the best possible solver for the type of problem addressed by the D-Wave architecture [9]. In another recent instance, an impressive large-scale simulation of spin dynamics on a quantum computer was reproduced by several groups using classical computers in a matter of weeks [10]. Therefore, the question if the problem has been compared to optimized classical solvers is always quite important to assess a quantum computing use case.
@@
On the more possitive note for quantum advantage, the possibility to find broadly applicable use cases for quantum computing can be substanially enhanced by investing efforts of transforming the original industry problem into a formulation appropriate for the quantum computer. As an example, solving differential equations is probably one of the key promising problems in applications of quantum computing. It has been shown that in principle we expect quantum computers to be able to solve differential equation with an exponential speedup [11, 12, 13]. But, a differential equation is never the actual industry problem. Usually, the actual problem requires the solutions of a differential equation. However, at this point only a small amount of work is published on the question how to transform the actual problem into the quantum computer without loosing the expected speedup. To the best of our knowledge, reference [14] is among the only studies addressing this topic directly in the context of manufacturing engineering and some work on problem-specific differential equations for machine learning [15, 16]. And much more should be possible.
In Table 1 we introduce our rating for these four ITBQ criteria and provide further details how to rank the use cases to track the process towards demonstrating quantum advantage. This table can be understood as a checklist and we will use the rating when describing the use cases in the following chapters. After discussing the two contrasting archetypes – Industrial Optimization and Abstract Spin Problems – in section II, we will discuss three problems at different levels of readiness to demonstrate quantum advantage: Nuclear Magnetic Resonance in sections III.1, Multireference Chemistry in section III.3 and Radicals in section III.4. Finally, we summarize our assessment for the discussed problems and conclude on the interpretation in section IV.
@@ II.1 Industrial Optimization: High Impact, Unproven Quantum Edge
Optimization problems underpin a vast array of real-world industrial processes. Mapping these problems onto quantum hardware is typically performed via transformation into the Quadratic Unconstrained Binary Optimization (QUBO) form—a mathematical structure well-suited to qubit-based quantum devices. However, this crucial Transform to Quantum step is not straightforward in practice. Industry problems are often high-dimensional and involve diverse constraints, which means they are neither quadratic nor unconstrained. Thus, the process of converting them to standardized QUBO instances is not straightforward and rarely automated. Consistently, robust and user-friendly software for this transformation is still in its infancy. Most transformations are tailored case-by-case, limiting the broader applicability of quantum optimization techniques.
@@ III.1.1 Identify industry problem
Nuclear magnetic resonance (NMR) spectroscopy is one of the most important analytical techniques in chemistry. When determining or elucidating the structure of organic compounds, it is usually NMR spectroscopy that provides the majority of information, complemented by other techniques such as infrared spectroscopy or mass spectrometry. The information content of NMR spectra is only superseded by X-ray crystallography. However, NMR spectra are usually recorded for substances in solution without needing to prepare crystals, rendering it much more suitable as a routine technique. In contrast to crystallographic structure determination, which often requires highly specialized training, interpretation of NMR spectra is considered a core skill and is introduced early in undergraduate chemistry cources [29]. Whenever organic (and many inorganic) substances are synthesized in a modern lab, it is likely that NMR spectra are recorded, too.
@@ III.1.2 Transform to quantum
III.2 Get the job done without the quantum computer
@@
III.2.1 Show quantum advantage
@@
However, when strong correlations extend across the system, as in the case of 1,2-di-tert-butyl-diphosphane (Fig. 8), the limitations of spin clustering become apparent. This molecule consists of 22 coupled spins, where cluster approximations cannot fully capture the spin interactions. Instead, solving the system exactly requires the exploitation of local SU(2) symmetry, which significantly reduces the computational complexity while maintaining full accuracy. Without symmetry considerations, the problem would require the diagonalization of a $2^{22}\times 2^{22}$-dimensional matrix, making classical exact diagonalization basically intractable. Given examples like 1,2-di-tert-butyl-diphosphane, it seams likely that parameters example and parameter regimes exist in which classical solutions remain challenging. At the same time it is clear that classical solvers can cover a large part of the relevant cases for NMR. Therefore we rate the probability to achieve quantum advantage with four stars (★★★★).
III.3 Multireference Chemistry
III.3.1 Identify industry problem
@@
Electronic structure calculations as an industry use case (with or without quantum computing) are tied to problems that can be solved with quantum chemistry in the first place: something that needs to be assessed on a case-by-case basis. We refer to ref. [69] for a discussion of examples with industrial relevance. Somewhat related is the problem of model building. A – comparably – simple system to simulate is a molecule in gas phase. Calculating the impact of a solvent is already more complicated: very often implicit solvation models are used, which surround the solute with a dielectric continuum to avoid simulating an ensemble of explicit solvent molecules. Setting up models of catalyst surfaces or of protein sites can be profoundly complicated tasks. An example is ref. [70], describing a combined multi-reference and multi-scale approach for a photosynthetic reaction center. Significant sources of quantitative errors in quantum chemistry include the computation of thermodynamic corrections [71], and relativistic effects become important for heavy elements. Thus, other aspects can be equally important as the electronic structure method itself.
@@ III.3.2 Transform to quantum
Dynamic correlation methods for smaller active spaces often employ internally contracted wave functions. Only reduced density matrices are needed as the output of the active space solver, making the methods independent of its specifics. Standard implementations of NEVPT2 or CASPT2 require four-particle density matrices. The number of tensor elements scales as $O\left(n_{\mathrm{act}}^{8}\right)$ with the number of active orbitals $n_{\mathrm{act}}$. More sophisticated dynamic correlation methods may require even higher order density matrices. This is reasonably unproblematic when the number of active orbitals is small. However, handling such tensors becomes unfeasible already in conventional calculations with large active spaces. Thus, the problem of recovering the dynamic multi-reference correlation energy in conjunction with conventional solvers for large active spaces remain without a universally valid solution. Suggested approaches are typically strongly tied to the specifics of the configuration interaction approximation employed, such as methods developed specifically for DMRG [98, 99] or for heat bath configuration interaction [100]. It seems likely that a method to recover dynamic correlation on top of a quantum computing solver would depend its specifics, too. We will discuss our own method, that allows a mapping of both active space and remaining dynamic correlation to quantum computers in the section III.4.
@@
III.3.3 Get the job done without the quantum computer
@@
It can be more complicated to assess the quality of the solution for the full system including its inactive orbitals, especially if dynamic correlation is accounted for. The variational principle does not apply to methods such as coupled cluster or NEVPT2, nor is it meaningful to compare to DFT total energies. Additionally, errors in electronic energies from conventional quantum chemistry methods are usually several orders of magnitude larger than errors in energy differences between geometries or electronic states. Therefore, a lower total energy does not imply that the method is superior in general: it may have worse error cancellation, leading to worse performance for computed properties. Emphasizing this point is important: total energies can be a useful criterion for approximations to the same method if the variational principle is adhered to, such as different full CI solver for identical active orbitals. However, comparing total energies of different methods, such as CASCI with NEVPT2, DFT or coupled cluster, will not lead to meaningful conclusions about their practical performance.
@@
III.3.4 Show quantum advantage
@@
One fundamental issue lies in the electronic structure of molecules at equilibrium. Many molecules in their ground states exhibit single-reference character, meaning their wavefunctions are dominated by a single Slater determinant. Classical methods, such as Hartree-Fock, coupled-cluster, or density functional theory (DFT), excel at handling single-reference systems, providing accurate ground-state energies with polynomial computational cost. Quantum computers, by contrast, are expected to shine in multireference scenarios, where strong electron correlation leads to wavefunctions with multiple significant determinants – a situation common in transition states, radicals (sell also section III.4), or certain transition metal complexes, but less prevalent in equilibrium geometries.
@@
III.4 Radicals
III.4.1 Identify industry problem
@@
There is also a significant academic relevance to studying diradicals because they play a fundamental role in understanding electronic correlation and reactivity in open-shell systems. Diradicals serve as classic examples for testing and developing new electronic structure theories because their electronic properties challenge the assumptions of traditional single-reference quantum chemistry methods. Additionally, diradicals are key intermediates in a variety of important chemical reactions, such as combustion, polymerization, and photochemical processes. Insights gained from research on diradicals enhance our understanding of spin dynamics, electron correlation effects, and non-trivial bonding situations, thereby advancing core concepts in theoretical and computational chemistry.
@@ III.4.2 Transform to quantum
After this procedure in which we used standard methods from quantum chemistry, we now preform transformations on the level of the Hamiltonian, to use the classically efficent static approximation to solve the random phase approximation, but also allow for a higher order approximation when mapping to a quantum computer. Since here we are considering diradicals, we separate the two orbitals hosting the two unpaired electrons and number them as the orbitals $1$ and $2$. We will call this the active space from now on. All remaining orbitals, both doubly occupied and empty, are numbered as $3,4,5,\dots,N$. These orbitals form the environment for the active space. After such separation, the Hamiltonian takes the form
@@
The Hamiltonian (14) can be seen as the final product of our Transform to Quantum efforts. From this point onward we can now try to solve the problem classically, which we will discuss in the next section, or we can try to map the problem to the quantum computer which will be discussed in section III.4.4. The derivation of the Hamiltonian has also been discussed in detail in Ref. [122].
@@
III.4.3 Get the job done without the quantum computer
@@
III.4.4 Show quantum advantage
@@ IV Conclusions
In the case of Radicals, the assessment considers both the relevance for chemistry and photomaterials and the characteristics of the current quantum chemistry approaches, which tend to be either computationally expensive or insufficiently robust. The given rating reflects the limited scope of the use case, as it targets a highly specific property, the singlet-triplet gap, of a particular class of compounds, namely diradicals. While this gap is an important parameter, it alone is not sufficient for the design of new molecules and materials. Pre- and post-processing has the main challenges being the assumption of the same geometric structure for singlet and triplet states, and the complex quantum chemistry computations needed to generate the input Hamiltonian. That said, it is broadly agreed that even achieving quantum advantage for purely academic research is a worthwhile goal. For benchmarking we note the availability of a broad range of conventional electronic structure methods that differ in computational cost, accuracy and robustness, alongside an efficient custom HQS-developed approach producing reasonable results at moderate costs. Quantum advantage has not yet been assessed is detail and remains to be explored further. But the specific transformations we have chosen allows for interesting possibilites when it comes to mapping the problem to the quantum computer.

### v2

@@ I Introduction
Quantum computing hardware has progressed substantially in the last few years. Several quantum computing hardware providers can relatively consistently achieve two qubit gates with more than 99% Fidelity. Under laboratory conditions, even 99.9% has been achieved in quite a few architectures, and at least one cloud-accessible device with these remarkable fidelities exists (Quantinuum). This all makes it even more important to drive use cases forward. A vast amount of work has been dedicated to testing various algorithms and approaches in the last few years [1, 2, 3, 4, 5, 6] and it is now all the more important to assess the state and quality of the existing ideas for use cases.
@@
However, it is often noticable that combining just these two criteria in a completely satisfactory way remains quite challenging. As we will discuss later, there is a vast number of studies on problems that are without any doubts of great industrial relevance, but lack substantiated arguments on how to achieve quantum advantage. Concurrently, it is highly nontrivial to connect algorithms with strictly provable quantum advantage to actual applications. Most prominently, this is the case for sampling experiments which are certainly of crucial importance as a fundamental performance benchmark, but until recently [7] lacked any connection to applications.
@@
The second criterion we intend to add is the need to invest as much effort to solve the given problem with classical computers as into the research on quantum computing. We call this criteria Get the Job Done Without the Quantum Computer. Of course, scientific research often has limited resources, so it might not always be possible to fully optimize classical methods before exploring the quantum ones. Nevertheless, at some stage in the process, it’s important to consider the status and availability of optimized classical solutions when judging any work on a quantum use case. As an example, one of the first claims advanced for an andvantage in using quantum computers was based on a benchmark performed by a D-Wave quantum computer [8] versus classical CPLEX solver, a standard tool for combinatorial optimization problems. However, this example did not stand up to scrutiny in the long run, since CPLEX is not the best possible solver for the type of problem addressed by the D-Wave architecture [9]. In another recent instance, an impressive large-scale simulation of spin dynamics on a quantum computer was reproduced by several groups using classical computers in a matter of weeks [10]. Therefore, the question if the problem has been compared to optimized classical solvers is always quite important to assess a quantum computing use case.
@@
On a more possitive note for quantum advantage, the possibility to find broadly applicable use cases for quantum computing can be substanially enhanced by investing efforts of transforming the original industry problem into a formulation appropriate for the quantum computer. As an example, solving differential equations is probably one of the key promising problems in applications of quantum computing. It has been shown that in principle we expect quantum computers to be able to solve differential equation with an exponential speedup [11, 12, 13]. But, a differential equation is never the actual industry problem. Usually, the actual problem requires the solution of a differential equation. However, at this point only a small amount of work is published on the question how to transform the actual problem onto the quantum computer without losing the expected speedup. To the best of our knowledge, reference [14] is among the only studies addressing this topic directly in the context of manufacturing engineering and some work on problem-specific differential equations for machine learning [15, 16]. And much more should be possible.
In Table 1 we introduce our rating for these four ITBQ criteria and provide further details how to rank the use cases to track the process towards demonstrating quantum advantage. This table can be understood as a checklist and we will use the rating when describing the use cases in the following chapters. After discussing the two contrasting archetypes – Industrial Optimization and Abstract Spin Problems – in section II, we will discuss three problems at different levels of readiness to demonstrate quantum advantage: Nuclear Magnetic Resonance in sections III.1, Multireference Chemistry in section III.2 and Radicals in section III.3. Finally, we summarize our assessment for the discussed problems and conclude on the interpretation in section IV.
@@ II.1 Industrial Optimization: High Impact, Unproven Quantum Edge
Optimization problems underpin a vast array of real-world industrial processes. Mapping these problems onto quantum hardware is typically performed via transformation into the Quadratic Unconstrained Binary Optimization (QUBO) form—a mathematical structure well-suited to qubit-based quantum devices. However, this crucial Transform to Quantum step is not straightforward in practice. Industry problems are often high-dimensional and involve diverse constraints, which means they are neither quadratic nor unconstrained. Thus, the process of converting them to standardized QUBO instances is not straightforward and rarely automated. Consequently, robust and user-friendly software for this transformation is still in its infancy. Most transformations are tailored case-by-case, limiting the broader applicability of quantum optimization techniques.
@@ III.1.1 Identify industry problem
Nuclear magnetic resonance (NMR) spectroscopy is one of the most important analytical techniques in chemistry. When determining or elucidating the structure of organic compounds, it is usually NMR spectroscopy that provides the majority of information, complemented by other techniques such as infrared spectroscopy or mass spectrometry. The information content of NMR spectra is only superseded by X-ray crystallography. However, NMR spectra are usually recorded for substances in solution without needing to prepare crystals, rendering it much more suitable as a routine technique. In contrast to crystallographic structure determination, which often requires highly specialized training, interpretation of NMR spectra is considered a core skill and is introduced early in undergraduate chemistry courses [29]. Whenever organic (and many inorganic) substances are synthesized in a modern lab, it is likely that NMR spectra are recorded, too.
@@ III.1.2 Transform to quantum
III.1.3 Get the job done without the quantum computer
@@
III.1.4 Show quantum advantage
@@
However, when strong correlations extend across the system, as in the case of 1,2-di-tert-butyl-diphosphane (Fig. 8), the limitations of spin clustering become apparent. This molecule consists of 22 coupled spins, where cluster approximations cannot fully capture the spin interactions. Instead, solving the system exactly requires the exploitation of local SU(2) symmetry, which significantly reduces the computational complexity while maintaining full accuracy. Without symmetry considerations, the problem would require the diagonalization of a $2^{22}\times 2^{22}$-dimensional matrix, making classical exact diagonalization basically intractable. Given examples like 1,2-di-tert-butyl-diphosphane, it seams likely that parameter examples and parameter regimes exist in which classical solutions remain challenging. At the same time it is clear that classical solvers can cover a large part of the relevant cases for NMR. Therefore we rate the probability to achieve quantum advantage with four stars (★★★★).
III.2 Multireference Chemistry
III.2.1 Identify industry problem
@@
Electronic structure calculations as an industry use case (with or without quantum computing) are tied to problems that can be solved with quantum chemistry in the first place: something that needs to be assessed on a case-by-case basis. We refer to ref. [69] for a discussion of examples with industrial relevance. This process necessarily involves model building to capture the essential physics and chemistry of the target system. For instance, a comparably simple system to simulate is a molecule in gas phase. Calculating the impact of a solvent is already more complicated: very often implicit solvation models are used, which surround the solute with a dielectric continuum to avoid simulating an ensemble of explicit solvent molecules. Setting up models of catalyst surfaces or of protein sites can be profoundly complicated tasks. An example is ref. [70], describing a combined multi-reference and multi-scale approach for a photosynthetic reaction center. Significant sources of quantitative errors in quantum chemistry include the computation of thermodynamic corrections [71], and relativistic effects become important for heavy elements. Thus, other aspects can be equally important as the electronic structure method itself.
@@ III.2.2 Transform to quantum
Dynamic correlation methods for smaller active spaces often employ internally contracted wave functions. Only reduced density matrices are needed as the output of the active space solver, making the methods independent of its specifics. Standard implementations of NEVPT2 or CASPT2 require four-particle density matrices. The number of tensor elements scales as $O\left(n_{\mathrm{act}}^{8}\right)$ with the number of active orbitals $n_{\mathrm{act}}$. More sophisticated dynamic correlation methods may require even higher order density matrices. This is reasonably unproblematic when the number of active orbitals is small. However, handling such tensors becomes unfeasible already in conventional calculations with large active spaces. Thus, the problem of recovering the dynamic multi-reference correlation energy in conjunction with conventional solvers for large active spaces remain without a universally valid solution. Suggested approaches are typically strongly tied to the specifics of the configuration interaction approximation employed, such as methods developed specifically for DMRG [98, 99] or for heat bath configuration interaction [100]. It seems likely that a method to recover dynamic correlation on top of a quantum computing solver would depend its specifics, too. We will discuss our own method, that allows a mapping of both active space and remaining dynamic correlation to quantum computers in the section III.3.
@@
III.2.3 Get the job done without the quantum computer
@@
It can be more complicated to assess the quality of the solution for the full system including its inactive orbitals, especially if dynamic correlation is accounted for. The variational principle does not apply to methods such as coupled cluster or NEVPT2, nor is it meaningful to compare to DFT total energies. Additionally, errors in electronic energies from conventional quantum chemistry methods are usually several orders of magnitude larger than errors in energy differences between geometries or electronic states. Therefore, a lower total energy does not imply that the method is superior in general: it may have worse error cancellation, leading to worse performance for computed properties. Emphasizing this point is important: total energies can be a useful criterion for approximations to the same method if the variational principle is adhered to, such as different full CI solvers for identical active orbitals. However, comparing total energies of different methods, such as CASCI with NEVPT2, DFT or coupled cluster, will not lead to meaningful conclusions about their practical performance.
@@
III.2.4 Show quantum advantage
@@
One fundamental issue lies in the electronic structure of molecules at equilibrium. Many molecules in their ground states exhibit single-reference character, meaning their wavefunctions are dominated by a single Slater determinant. Classical methods, such as Hartree-Fock, coupled-cluster, or density functional theory (DFT), excel at handling single-reference systems, providing accurate ground-state energies with polynomial computational cost. Quantum computers, by contrast, are expected to shine in multireference scenarios, where strong electron correlation leads to wavefunctions with multiple significant determinants – a situation common in transition states, radicals (sell also section III.3), or certain transition metal complexes, but less prevalent in equilibrium geometries.
@@
III.3 Radicals
III.3.1 Identify industry problem
@@
There is also a significant academic relevance to studying diradicals because they play a fundamental role in understanding electronic correlation and reactivity in open-shell systems. Diradicals serve as classic examples for testing and developing new electronic structure theories because their electronic properties challenge the assumptions of traditional single-reference quantum chemistry methods. Additionally, diradicals are key intermediates in a variety of important chemical reactions, such as combustion, polymerization, and photochemical processes. Insights gained from research on diradicals enhance our understanding of spin dynamics, electron correlation effects, and nontrivial bonding situations, thereby advancing core concepts in theoretical and computational chemistry.
@@ III.3.2 Transform to quantum
After this procedure in which we used standard methods from quantum chemistry, we now perform transformations on the level of the Hamiltonian, to use the classically efficent static approximation to solve the random phase approximation, but also allow for a higher order approximation when mapping to a quantum computer. Since here we are considering diradicals, we separate the two orbitals hosting the two unpaired electrons and number them as the orbitals $1$ and $2$. We will call this the active space from now on. All remaining orbitals, both doubly occupied and empty, are numbered as $3,4,5,\dots,N$. These orbitals form the environment for the active space. After such separation, the Hamiltonian takes the form
@@
The Hamiltonian (14) can be seen as the final product of our Transform to Quantum efforts. From this point onward we can now try to solve the problem classically, which we will discuss in the next section, or we can try to map the problem to the quantum computer which will be discussed in section III.3.4. The derivation of the Hamiltonian has also been discussed in detail in Ref. [122].
@@
III.3.3 Get the job done without the quantum computer
@@
III.3.4 Show quantum advantage
@@ IV Conclusions
In the case of Radicals, the assessment considers both the relevance for chemistry and photomaterials and the characteristics of the current quantum chemistry approaches, which tend to be either computationally expensive or insufficiently robust. The given rating reflects the limited scope of the use case, as it targets a highly specific property, the singlet-triplet gap, of a particular class of compounds, namely diradicals. While this gap is an important parameter, it alone is not sufficient for the design of new molecules and materials. Pre- and post-processing has the main challenges being the assumption of the same geometric structure for singlet and triplet states, and the complex quantum chemistry computations needed to generate the input Hamiltonian. That said, it is broadly agreed that even achieving quantum advantage for purely academic research is a worthwhile goal. For benchmarking we note the availability of a broad range of conventional electronic structure methods that differ in computational cost, accuracy and robustness, alongside an efficient custom HQS-developed approach producing reasonable results at moderate costs. Quantum advantage has not yet been assessed in detail and remains to be explored further. But the specific transformations we have chosen allows for interesting possibilites when it comes to mapping the problem to the quantum computer.